### v1

# Non-Asymptotic Performance Analysis of DOA Estimation Based on Real-Valued Root-MUSIC

Abstract
This paper presents a systematic theoretical performance analysis of the Real-Valued root-MUSIC (RV-root-MUSIC) algorithm under non-asymptotic conditions. However, RV-root-MUSIC suffers from the problem of estimation ambiguity for the mirror roots, therefore the conventional beamforming (CBF) technique is typically employed to filter out the mirror roots. Through the equivalent subspace based on the conjugate extension method and the equivalence of perturbations for both true roots and mirror roots , this paper provides a comprehensive investigation of three critical aspects: noise subspace perturbation, true root perturbation, and mirror root perturbation characteristics in the RV-root-MUSIC algorithm. The statistical model is established and the generalized expression of perturbation is further developed. The simulation results show the correctness and validity of the derived statistical characteristics. The results provide a solid theoretical foundation for optimizing the parameter selection of DOA estimation in practical applications, particularly in radar systems, communication networks, and intelligent sensing technologies.
Index Terms—  RV-root-MUSIC, Direction-of-Arrival (DOA) estimation, non-asymptotic performance analysis, perturbation-theoretic expressions
1 Introduction
Direction of Arrival (DOA) estimation is one of nonlinear parameter estimation problems. Within this domain, the accurate acquisition of target parameters remains a central challenge in signal processing research. High-precision parameter estimation serves as a critical enabler for target perception across diverse applications, including radar detection, wireless communications, sonar localization, and medical imaging[1, 2, 3, 4, 5]. However, inherent noise contamination during signal transmission and reception introduces unavoidable errors. Consequently, estimation accuracy emerges as a key metric for evaluating algorithmic performance. To develop robust error quantification frameworks and systematic performance analysis methodologies is significant for DOA estimation [6].
The Multiple Signal Classification (MUSIC)[7] algorithm has revolutionized DOA estimation by achieving super-resolution performance. However, its high computational complexity due to spectral peak search limits its practical applications. To address this, the root-MUSIC[8] algorithm was developed, significantly reducing computational load. Despite these advancements, MUSIC and its variants operate in the complex domain, which still poses efficiency challenges[9]. Consequently, real-valued subspace-based algorithms, such as Real-Valued MUSIC (RV-MUSIC) [10] and Real-Valued root-MUSIC (RV-root-MUSIC) [11, 12], have been proposed to further enhance computational efficiency. While the conventional root-MUSIC exhibits satisfactory performance under high signal-to-noise ratio (SNR) and asymptotic conditions, its estimation accuracy degrades significantly in low-SNR scenarios or limited snapshots (non-asymptotic conditions). Performance analysis of conventional complex-domain DOA estimation algorithms has been conducted[13, 14], while the performance of real-valued algorithms remains to be analysed.
This paper presents a non-asymptotic performance analysis of RV-root-MUSIC. Theoretical error expressions are derived for both true roots and mirror roots, with their validity confirmed through Monte Carlo simulations demonstrating agreement between theory and simulation.
2 PRELIMINARIES
2.1 The model of the signal and arrays
Considering $K$ far-field narrowband signals arriving on uniform linear arrays (ULA) with $L$ elements, the ideal received array signal output at $M$ snapshots can be expressed as:
$\bm{X}(t)=\bm{A}(\theta)\bm{S}(t)$ (1)
where $\bm{X}(t)\in\mathbb{C}^{L\times M}$ is the ideal received signal matrix, $\bm{A}(\theta)\in\mathbb{C}^{L\times K}$ is the array manifold matrix of ULA, consisting of steering vectors as its columns, expressed as:
$\bm{a}(\theta_{k})=\begin{bmatrix}1&e^{-j\frac{2\pi d}{\lambda}sin\theta_{k}}&\cdots&e^{-j(L-1)\frac{2\pi d}{\lambda}sin\theta_{k}}\\ \end{bmatrix}^{T}$ (2)
where $d$ denotes the inter-element spacing, and $\lambda$ represents the signal wavelength. $\bm{S}(t)\in\mathbb{C}^{K\times M}$ is the source signal matrix.
When the perturbation $\Delta\bm{X}$ is taken into account, the above expression can be rewritten as:
$\bm{\tilde{X}}(t)=\bm{X}(t)+\Delta\bm{X}=\bm{A}(\theta)\bm{S}(t)+\bm{N}(t)$ (3)
where $\bm{N}(t)\in\mathbb{C}^{L\times M}$ is the additive noise matrix, typically modeled as complex circularly symmetric Gaussian.
In DOA estimation, the covariance matrix of received signals $\bm{\tilde{R}}_{xx}$ is commonly employed instead of the raw signal matrix $\bm{\tilde{X}}$ for analysis. This approach offers advantages such as accelerated computation through eigenvalue decomposition (EVD) and dimensionality reduction. $\bm{\tilde{R}}_{xx}$ is formulated as:
$\bm{\tilde{R}}_{xx}=\frac{1}{M}\sum_{t=1}^{M}\bm{\tilde{X}}(t)\bm{\tilde{X}}^{H}(t)$ (4)
2.2 Review about RV-root-MUSIC
The RV-root-MUSIC algorithm takes the real part of the covariance matrix, where the complex calculation of original root-MUSIC algorithm can be converted into the efficient real-valued one. Then EVD can be performed to extract the signal and noise subspace:
$\bm{\tilde{R}}_{real}=\mathrm{Re}(\bm{\tilde{R}}_{xx})=\bm{\tilde{\mathbb{E}}}_{s}\bm{\tilde{\Lambda}}_{s}\bm{\tilde{\mathbb{E}}}^{T}_{s}+\bm{\tilde{\mathbb{E}}}_{n}\bm{\tilde{\Lambda}}_{n}\bm{\tilde{\mathbb{E}}}^{T}_{n}$ (5)
where $span(\bm{\mathbb{E}}_{s})$ represents the real-valued eigenvectors spanning the signal subspace, with $\bm{\Lambda}_{s}$ denoting its associated diagonal eigenvalue matrix. The noise subspace components $span(\bm{\mathbb{E}}_{n})$ and $\bm{\Lambda}_{n}$ are defined analogously using real-valued quantities and $\bm{\tilde{R}}_{real}=\bm{R}_{real}+\Delta\bm{R}_{real}$, $\bm{\tilde{\mathbb{E}}}_{s}=\bm{\mathbb{E}}_{s}+\Delta\bm{\mathbb{E}}_{s}$, $\bm{\tilde{\Lambda}}_{s}=\bm{\Lambda}_{s}+\Delta\bm{\Lambda}_{s}$, $\bm{\tilde{\mathbb{E}}}_{n}=\bm{\mathbb{E}}_{n}+\Delta\bm{\mathbb{E}}_{n}$, $\bm{\tilde{\Lambda}}_{n}=\bm{\Lambda}_{n}+\Delta\bm{\Lambda}_{n}=\Delta\bm{\Lambda}_{n}$.
Based on the defined noise subspace, the root-finding polynomial for DOA estimation algorithm can be constructed as:
$f_{RV-root-MUSIC}(z)=\bm{p}^{T}(z^{-1})\bm{\mathbb{E}}_{n}\bm{\mathbb{E}}^{T}_{n}\bm{p}(z)$ (6)
where $\bm{p}(z)$ is defined as $1,z,z^{2},\cdots,z^{M-1}]$, $z=e^{jw}$, and $w=\frac{2\pi dsin\theta}{\lambda}$. Then based on the root-solving results, the DOA estimates can be calculated through:
$\theta=arcsin[\frac{\lambda}{2\pi d}arg(z)]$ (7)
As demonstrated in [11], RV-root-MUSIC suffers from the mirror roots, where the relationship $\theta_{mirror}=-\theta_{true}$ is satisfied. Thus the conventional beamforming (CBF) technique is typically employed to filter out the mirror roots.
3 Non-asymptotic Performance Analysis of RV-root-MUSIC
This paper makes a statistical characteristic analysis of real roots and mirror roots about RV-root-MUSIC under non-progressive conditions and also verifies the estimation accuracy with respect to the number of array elements.
The polynomial root-finding function of RV-root-MUSIC can be rewritten as:
$f=AK(z)G(z)$ (8)
where $K(z)$ denotes the roots related to the DOA roots:
$K(z)=(1-r_{k}z^{-1})(1-r_{k}^{*}z)(1-\frac{1}{r_{k}}z^{-1})(1-\frac{1}{r_{k}^{*}}z)$ (9)
while $G(z)$ represents other extraneous roots:
$G(z)=\prod\limits_{j=1,j\neq k}^{\left\lfloor\frac{L-1}{2}\right\rfloor}(1-r_{j}\frac{1}{z})(1-r_{j}^{*}z)(1-\frac{1}{r_{j}}\frac{1}{z})(1-\frac{1}{r_{j}^{*}}z)$ (10)
As mentioned, RV-root-MUSIC can provide the mirror roots, therefore we can further divide $K(z)$ into the real root parts $K_{true}(z)=(1-r_{k}z^{-1})(1-r_{k}^{*}z)$ and the mirror root parts $K_{mirror}(z)=(1-\frac{1}{r_{k}}z^{-1})(1-\frac{1}{r_{k}^{*}}z)$:
$K(z)=K_{true}(z)K_{mirror}(z)=K_{t}(z)K_{m}(z)$ (11)
To derive the expression for the DOA estimation bias, we first define a logarithmic function $f(x)=lnx$. Assume that $x=r_{k}$ is the root corresponding to the $k$-th DOA, and $\Delta r_{k}$ denotes the error in estimating this root. The first-order Taylor expansion can be performed on $f(x)$ to obtain:
$f(\tilde{r}_{k})=lnc+j\tilde{\omega}_{k}=lnc+j\omega+f^{\prime}(r_{k})\Delta r_{k}$ (12)
where $\tilde{r}_{k}=ce^{j\tilde{\omega}_{k}}=r_{k}+\Delta r_{k}$, $\tilde{\omega}_{k}=\omega_{k}+\Delta\omega_{k}$, and $c$ is the amplitude perturbation caused by additive noise.
Using $f^{\prime}(r_{k})=1/r_{k}$ and (12) can be simplified by canceling common terms, then take the imaginary part of both sides, we obtain:
$\Delta\omega_{k}=\mathrm{Im}(\frac{\Delta r_{k}}{r_{k}})$ (13)
Note that when $\Delta\theta_{k}\to 0$, the relationship $\Delta\omega_{k}=\frac{2\pi d}{\lambda}\cos\theta_{k}\Delta\theta_{k}$ is satisfied, we can further derive the bias expression of the final DOA estimation:
$\Delta\theta_{k}=C_{k}\,\mathrm{Im}(\frac{\Delta r_{k}}{r_{k}})$ (14)
where the scaling factor is $C_{k}=\frac{\lambda}{2\pi d\cos\theta_{k}}$. The derivation for mirror DOA follows the same procedure.
3.1 Deviation of the true DOAs
To derive the perturbation expression, $\mathrm{Im}(\frac{\Delta r_{k}}{r_{k}})$ in (14) is necessary to be computed. The derivation can be decomposed into two steps:
step 1. Take the partial derivative with respect to the root $z$, and substitute $z=r_{k}$. Considering the perturbation of $r$, replace $r_{k}$ with $\tilde{r}=r_{k}+\Delta(r_{k})$ and only retain the first-order term to obtain the perturbation expression of $r_{k}$, as below:
$\begin{split}\left.\frac{\partial f(z,\tilde{r})}{\partial z}\right|_{z=r_{k}}&=A[K_{t}(r_{k})^{{}^{\prime}}[K_{m}(r_{k})G(r_{k})]\\ &\quad\left.+K(r_{k})[K_{m}(r_{k})G(r_{k})]^{{}^{\prime}}]\right|_{\tilde{r}=r_{k}+\Delta r_{k}}\\ &=A[K_{t}(r_{k})^{{}^{\prime}}K_{m}(r_{k})G(r_{k})]\\ &\doteq 2jAr_{k}^{*}Im(\frac{\Delta r_{k}}{r_{k}})K_{m}(r_{k})G(r_{k})\end{split}$ (15)
step 2. As in step 1, but now consider the perturbation of $\bm{\mathbb{E}}_{n}$, replace $\bm{\mathbb{E}}_{n}$ with $\bm{\tilde{\mathbb{E}}}_{n}=\bm{\mathbb{E}}_{n}+\Delta\bm{\mathbb{E}}_{n}$ and retain only the first-order terms, we obtain:
[…]
where $p^{(1)}(z)=\frac{\partial p(z)}{\partial z}$. Substitute $\bm{\tilde{\mathbb{E}}}_{n}=\bm{\mathbb{E}}_{n}+\Delta\bm{\mathbb{E}}_{n}$ and recognize the relation below:
[…]
Given that $\bm{\mathbb{E}}^{T}_{n}\bm{p}(r_{k})=0$ and the relation of (17), (16) can be simplified and we can further obtain the perturbation expression of $\bm{\mathbb{E}}_{n}$:
$\left.\frac{\partial f(z,\bm{\tilde{\mathbb{E}}}_{n})}{\partial z}\right|_{z=r_{k}}\doteq 2jr_{k}^{*}\mathrm{Im}[r_{k}\bm{p}^{H}(r_{k})\Delta\bm{\mathbb{E}}_{n}\bm{\mathbb{E}}_{n}^{T}\bm{p}^{(1)}(r_{k})]\\$ (18)
Since the perturbation of $r_{k}$ originates from the perturbation of $\bm{\mathbb{E}}_{n}$, and both ultimately stem from noise, we can equate (15) and (18) to obtain:
$\mathrm{Im}(\frac{\Delta r_{k}}{r_{k}})=\frac{\mathrm{Im}[r_{k}\bm{p}^{H}(r_{k})\Delta\bm{\mathbb{E}}_{n}\bm{\mathbb{E}}_{n}^{T}\bm{p}^{(1)}(r_{k})]}{AK_{m}(r_{k})G(r_{k})}$ (19)
Substituting to the derived equation (14) above yields:
$\Delta\theta_{k}=\frac{\lambda}{2\pi dcos\theta_{k}}\frac{\mathrm{Im}[r_{k}\bm{p}^{H}(r_{k})\Delta\bm{\mathbb{E}}_{n}\bm{\mathbb{E}}_{n}^{T}\bm{p}^{(1)}(r_{k})]}{AK_{m}(r_{k})G(r_{k})}$ (20)
3.2 Deviation of the mirror DOAs
Since the relationship $r_{\varphi_{k}}=r_{\theta_{k}}^{*}$ is satisfied for the mirror roots $\varphi_{k}=-\theta_{k}$, their analysis can be directly obtained by substituting $z=r_{k}^{*}$ into the derivations from the previous section.
Following the same procedure as in the previous step 1, we can derive the perturbation expression of $r_{k}^{*}$:
$\left.\frac{\partial f(z,\tilde{r})}{\partial z}\right|_{z=r_{k}^{*}}\doteq 2jAr_{k}\mathrm{Im}(\frac{\Delta r_{k}^{*}}{r_{k}^{*}})K_{t}(r_{k}^{*})G(r_{k}^{*})$ (21)
and the perturbation expression of $\bm{\mathbb{E}}_{n}$ as step 2:
$\left.\frac{\partial f(z,\bm{\tilde{\mathbb{E}}}_{n})}{\partial z}\right|_{z=r_{k}^{*}}\doteq 2jr_{k}\mathrm{Im}[r_{k}^{*}\bm{p}^{H}(r_{k}^{*})\Delta\bm{\mathbb{E}}_{n}\bm{\mathbb{E}}_{n}^{T}\bm{p}^{(1)}(r_{k}^{*})]\\$ (22)
By the equivalence of perturbations, we can get:
$\mathrm{Im}(\frac{\Delta r_{k}^{*}}{r_{k}^{*}})=\frac{\mathrm{Im}[r_{k}^{*}\bm{p}^{H}(r_{k}^{*})\Delta\bm{\mathbb{E}}_{n}\bm{\mathbb{E}}_{n}^{T}\bm{p}^{(1)}(r_{k}^{*})]}{AK_{t}(r_{k}^{*})G(r_{k}^{*})}$ (23)
and further obtain:
$\Delta\varphi_{k}=\frac{\lambda}{2\pi dcos\varphi_{k}}\frac{\mathrm{Im}[r_{k}^{*}\bm{p}^{H}(r_{k}^{*})\Delta\bm{\mathbb{E}}_{n}\bm{\mathbb{E}}_{n}^{T}\bm{p}^{(1)}(r_{k}^{*})]}{AK_{t}(r_{k}^{*})G(r_{k}^{*})}$ (24)
It should be noted that $\Delta r_{k}^{*}$ represents the perturbation at $z=r_{k}^{*}$, while $(\Delta r_{k})^{*}$ denotes the complex conjugate of the perturbation at $z=r_{k}$. These are fundamentally different quantities, as: $\Delta r_{k}^{*}\neq(\Delta r_{k})^{*}$, $(\Delta r_{k}^{*})^{*}\neq\Delta r_{k}$. This reflects how the function’s perturbation behavior varies at different evaluation points.
3.3 Deviation of the Noise subspace
Both (20) and (24) reveal that biases $\Delta\theta_{k}$ and $\Delta\varphi_{k}$ originate from perturbations in the noise subspace. The noise subspace perturbation is derived as follow:
We define a new signal model $\bm{X}_{v}=[\bm{X}\hskip 5.0pt\bm{X}^{*}]$, where the following properties are satisfied after considering the noise:
$\begin{split}\bm{\tilde{X}}_{v}\bm{\tilde{X}}_{v}^{H}=2\mathrm{Re}(\bm{\tilde{X}}\bm{\tilde{X}}^{H})=2M\mathrm{Re}(\bm{\tilde{R}}_{xx})\end{split}$ (25)
where $\bm{\tilde{X}}_{v}=\bm{X}_{v}+\Delta\bm{X}_{v}$, $\Delta\bm{X}_{v}=[\Delta\bm{X}\hskip 5.0pt\Delta\bm{X}^{*}]$. Then Singular Value Decomposition (SVD) of $\bm{\tilde{X}}_{v}$ and EVD of $\bm{\tilde{R}}_{real}$ yield the following results:
[…]
where $\bm{\tilde{U}}_{s}=\bm{U}_{s}+\Delta\bm{U}_{s}$, $\bm{\tilde{W}}_{s}=\bm{W}_{s}+\Delta\bm{W}_{s}$, $\bm{\tilde{V}}_{s}=\bm{V}_{s}+\Delta\bm{V}_{s}$, $\bm{\tilde{U}}_{n}=\bm{U}_{n}+\Delta\bm{U}_{n}$, $\bm{\tilde{W}}_{n}=\bm{W}_{n}+\Delta\bm{W}_{n}$, $\bm{\tilde{V}}_{n}=\bm{V}_{n}+\Delta\bm{V}_{n}$.
Substituting (26) into (25) can yield:
$\bm{\tilde{U}}_{s}\bm{\tilde{W}}_{s}^{2}\bm{\tilde{U}}_{s}^{H}+\bm{\tilde{U}}_{n}\bm{\tilde{W}}_{n}\bm{\tilde{W}}_{n}^{T}\bm{\tilde{U}}_{n}^{H}=2M(\bm{\tilde{\mathbb{E}}}_{s}\bm{\tilde{\Lambda}}_{s}\bm{\tilde{\mathbb{E}}}^{T}_{s}+\bm{\tilde{\mathbb{E}}}_{n}\bm{\tilde{\Lambda}}_{n}\bm{\tilde{\mathbb{E}}}^{T}_{n})$ (27)
Observing (27), we can obtain following relationships:
$\begin{split}\bm{\tilde{U}}_{s}=\bm{\tilde{\mathbb{E}}}_{s},\bm{\tilde{U}}_{n}=\bm{\tilde{\mathbb{E}}}_{n},\bm{\tilde{W}}_{s}^{2}=2M\bm{\tilde{\Lambda}}_{s},\bm{\tilde{W}}_{n}\bm{\tilde{W}}_{n}^{T}=2M\bm{\tilde{\Lambda}}_{n}\end{split}$ (28)
As derived in [13], for ULA, the perturbation of the noise subspace in the singular value domain of the received signal matrix $\bm{\tilde{X}}$ can be expressed as:
$\Delta\bm{U}_{o}=-\bm{U}_{s}\bm{W}_{s}^{-1}\bm{V}_{s}^{H}\Delta\bm{X}^{H}\bm{U}_{o}$ (29)
Based on (29) and the equivalent noise subspace characterization, we derive the noise subspace perturbation for RV-root-MUSIC as:
$\Delta\bm{\mathbb{E}}_{n}=\Delta\bm{U}_{n}=-\bm{U}_{s}\bm{W}_{s}^{-1}\bm{V}_{s}^{H}\Delta\bm{X}_{v}^{H}\bm{U}_{n}$ (30)
where $\Delta\bm{X}_{v}=\bm{N}_{v}=[\Delta\bm{X}\hskip 3.99994pt\Delta{\bm{X}}^{*}]=[\bm{N}\hskip 5.0pt\bm{N}^{*}]$.
3.4 Root Pairs of RV-root-MUSIC in even array elements
Theorem: For ULA with even elements, at least one root pair degenerates onto the real axis, corresponding to spurious DOA at $0^{\circ}$ or $180^{\circ}$ that align with the ULA axis. (see Fig. 1)
(a) 8 elements
(b) 9 elements
Fig. 1: Roots of RV-root-MUSIC in varying array elements
Proof: An odd-degree polynomial can be expressed as:
$y=a_{n}x^{n}+a_{n-1}x^{n-1}+\textperiodcentered\textperiodcentered\textperiodcentered\textperiodcentered\textperiodcentered\textperiodcentered+a_{1}x+a_{0}$ (31)
where $a_{n}\neq 0$ and $n$ is a positive odd integer. Dividing both sides by $x^{n}$ and taking the limit as $x\to\infty$ yields:
$\lim_{x\to\infty}\frac{y}{x^{n}}=\lim_{x\to\infty}a_{n}+\frac{a_{n-1}}{x}+\cdots=a_{n}$ (32)
Furthermore, since $n$ is odd, we have:
$\begin{split}\lim_{x\to+\infty}x^{n}&=+\infty\quad\lim_{x\to-\infty}x^{n}=$-$\infty\end{split}$ (33)
By the continuity of $y$ and its behavior at infinity, it can be obtained that there exist points in $\Re$ where $y$ changes sign. The Intermediate Value Theorem guarantees the existence of at least one real root satisfying $y=0$. Extending to the complex domain, for even-numbered arrays, (6) will transform into an odd-degree polynomial, and there must exist a pair of reciprocal roots on the real axis. Proof completed.
Therefore, for the framework proposed in (10), we derive a correction factor $R(a)$ for even-numbered arrays as follows:
$R(a)=a^{2}-2aRe(r_{k})+1=a^{2}-2aRe(r_{k}^{*})+1$ (34)
where $a$ is the real root previously discussed. To get more precise perturbation value, just multiply (20) and (24) by $\frac{1}{R(a)}$.
4 Statistical Properties of True and Mirror DOA Perturbations
Substituting (30) into (20) and (24), identical forms can be observed. To simplify performance analysis model, the general perturbation expression $\Delta_{k}$ can be expressed as:
$\Delta_{k}=-\frac{\mathrm{Im}(\bm{\beta}_{k}^{H}\bm{N}_{v}^{H}\bm{\alpha}_{k})}{\gamma_{k}}$ (35)
where the parameters $\bm{\alpha}_{k}$, $\bm{\beta}_{k}$, $\gamma_{k}$ are specified in Table 1.
Given that bias has zero mean and each element of the random variable $\Delta\bm{X}_{v}$ has variance $\sigma_{n}^{2}/2$ for both real and imaginary components, the MSE of bias is as follows:
$E_{\bm{N}_{v}}(\Delta_{k})^{2}=\frac{E_{\bm{N}_{v}}[\left|\bm{\beta}_{k}^{H}\bm{N}_{v}^{H}\bm{\alpha}_{k}\right|^{2}]}{2\gamma_{k}^{2}}=\frac{\left\|\bm{\alpha}_{k}\right\|^{2}\left\|\bm{\beta}_{k}\right\|^{2}\sigma_{n}^{2}}{2\gamma_{k}^{2}}$ (36)
Table 1: The parameters in the general expression
$\bm{\alpha}_{k}$ $\bm{\beta}_{k}$ $\gamma_{k}$
$\Delta\theta_{k}$ $\bm{\mathbb{E}}_{n}\bm{\mathbb{E}}_{n}^{T}\bm{p}^{(1)}(r_{k})C_{k}r_{k}$ $\bm{V}_{s}\bm{W}_{s}^{-1}\bm{U}_{s}^{H}\bm{p}(r_{k})$ $AK_{m}(r_{k})G(r_{k})$
$\Delta\varphi_{k}$ $\bm{\mathbb{E}}_{n}\bm{\mathbb{E}}_{n}^{T}\bm{p}^{(1)}(r_{k}^{*})C_{k}r_{k}^{*}$ $\bm{V}_{s}\bm{W}_{s}^{-1}\bm{U}_{s}^{H}\bm{p}(r_{k}^{*})$ $AK_{t}(r_{k}^{*})G(r_{k}^{*})$
5 Simulation Experiments and Results
To analyze the statistical impact of noise on DOA estimation, we conduct Monte Carlo simulations to examine the variation of estimation bias under zero-mean additive white Gaussian noise conditions.
A $9$-element ULA with $d=\lambda/2$ is used in the simulation, with two uncorrelated sources at $30^{\circ}$ and $50^{\circ}$. DOA estimation RMSEs for the source at $30^{\circ}$ from $1000$ Monte Carlo trials are investigated under $2$ conditions:
condition 1: Varying the SNR from $0dB$ to $20dB$ in $2dB$ steps with $200$ snapshot, and the result is shown in Fig. 2.
condition 2: Varying the snapshots $2^{n}$, where $n$ varies from $5$ to $12$ in $1$ step at $10dB$ SNR, with the result shown in Fig. 3.
(a) True DOA
(b) Mirror DOA
Fig. 2: The trend of estimation RMSE in SNR variation
From Fig. 2 and Fig. 3, it can be seen that the simulation performance closely aligns with the theoretical performance.
(a) True DOA
(b) Mirror DOA
Fig. 3: The trend of estimation RMSE in Snapshots variation
6 Summary
This paper formulated a framework of the non-asymptotic performance analysis of the RV-root-MUSIC algorithm, deeply analyzed the perturbation characteristics of the noise subspace, the true roots and the mirror roots, established a statistical model, and conducted simulation experiments. The final results show that the research in this paper can provide a tighter theoretical lower bound for the performance of the RV-root-MUSIC algorithm. In addition, the study on the parity of the array elements in this paper provides guidance for the performance improvement of the algorithm: that is, choosing an odd number of array elements to ensure the pairing phenomenon of the roots and thereby enhance the robustness of the algorithm.
7 Acknowledgment
The work of X. T. Meng is supported by the Shandong Provincial Natural Science Foundation (No.ZR2024QF068) and the Aeronautical Science Foundation of China (No.2023Z037077002). The work of F. Gini and M. Greco has been partially supported by the Italian Ministry of Education and Research (MUR) in the framework of the FoReLab project (Departments of Excellence).